# The CFT of SLE loop measures and the Kontsevich–Suhov conjecture

### v1

@@ 6. Uniqueness of restriction measures
The existence part is already known [Zha21], so we only need to show uniqueness. Let $\tilde{\nu}$ be a measure on the space of Jordan curves in the Riemann sphere satisfying the assumptions of Theorem 1.9. By conformal invariance, $\tilde{\nu}$ is completely characterised by its restriction to $\mathcal{J}_{0,\infty}$ (which we keep denoting $\tilde{\nu}$). Conformal invariance also implies that (up to rescaling by a multiplicative constant) $\tilde{\nu}=\pi^{-1/2}\mathrm{d}c\otimes\tilde{\nu}^{\#}$ for some probability measure $\tilde{\nu}^{\#}$ on $\mathcal{E}$. Let $\tilde{\mathbb{E}}^{\#}$ be the expectation w.r.t. $\tilde{\nu}^{\#}$. By Möbius invariance $z\mapsto 1/z$, $\Theta$ is self-adjoint, so $\tilde{\mathcal{Q}}(F,G):=\langle F,\Theta G\rangle_{L^{2}(\tilde{\nu})}$ is a Hermitian form on $L^{2}(\tilde{\nu})$.
By definition, the family of measures $\tilde{\nu}_{D}:=\mathds{1}_{\eta\subset D}e^{\frac{c_{\mathrm{m}}}{2}\Lambda^{*}(\eta,\partial D)}\mathrm{d}\tilde{\nu}(\eta)$ indexed by simply connected domains $D$ satisfies the following restriction property:
$\mathrm{d}\tilde{\nu}_{D_{2}}(\eta)=\mathds{1}_{\eta\subset D_{2}}e^{\frac{c_{\mathrm{m}}}{2}\Lambda_{D_{1}}(\eta,D_{2}^{c})}\mathrm{d}\tilde{\nu}_{D_{1}}(\eta),$
for all simply connected domains $D_{1},D_{2}$ with $\bar{D}_{2}\subset D_{1}$. Now, the conclusion of Proposition 4.5 remains true if we replace $\nu_{\mathbb{D}}$ by $\tilde{\nu}_{\mathbb{D}}$ (the proof only uses the restriction property), so $\tilde{\nu}_{\mathbb{D}}$ satisfies the integration by parts formula of Proposition 4.5. To get an integration by parts in the sphere, we proceed as before by noting that $|\log\epsilon|^{\frac{c_{\mathrm{m}}}{2}}\tilde{\nu}_{\epsilon\mathbb{D}^{*}}\to\tilde{\nu}$ weakly as $\epsilon\to 0$. Indeed, by the restriction property of $\tilde{\nu}$,
@@
and $\Lambda^{*}(\eta,\epsilon\mathbb{S}^{1})=-\log|\log\epsilon|+o(1)$ (this term does not depend on the specific choice of MKS measure and is thus the same as the one in [Zha21, page 385]). It follows that $\tilde{\nu}$ also satisfies the integration by parts formula of Theorem 4.1, i.e. we have the relations $\mathbf{L}_{n}^{*}=\Theta\circ\mathbf{L}_{-n}\circ\Theta$ and $\bar{\mathbf{L}}_{n}^{*}=\Theta\circ\bar{\mathbf{L}}_{-n}\circ\Theta$ for all $n\in\mathbb{Z}$, where the adjoint is w.r.t. to the $L^{2}(\tilde{\nu})$-inner product. In particular, $\tilde{\mathcal{Q}}$ implements the Shapovalov form of the representations and it satisfies the conclusions of Theorem 1.7 with $\tilde{\mathrm{R}}(\lambda):=\tilde{\mathbb{E}}^{\#}[e^{-2\lambda\tilde{c}_{0}}]$ in place of $\mathrm{R}(\lambda)$.
@@ 7.7. Couplings with the Gaussian free field
Following Sheffield’s seminal article [She16], many couplings have been proved between SLE and the Gaussian free field (GFF). This includes a recent proof that the independent pair of an SLE loop with a quantum sphere (roughly speaking, the GFF on the sphere) can be obtained as the conformal welding of two independent quantum discs conditioned on their quantum length [AHS21].

### v2

@@ 6. Uniqueness of restriction measures
The existence part is already known [Zha21], so we only need to show uniqueness. A conformal invariant Borel measure on the space $\mathcal{J}$ is completely characterised by its restriction to $\mathcal{J}_{0,\infty}$. Since the restriction to $\mathcal{J}_{0,\infty}$ of a restriction measure on $\mathcal{J}$ (in the sense of Definition 1.8) is a weak restriction measure on $\mathcal{J}_{0,\infty}$, it is enough to prove uniqueness of weak restriction measures on $\mathcal{J}_{0,\infty}$. Let $\tilde{\nu}$ be such a Borel measure on $\mathcal{J}_{0,\infty}$. We first claim that $\tilde{\nu}$ is scale invariant, i.e. $\tilde{\nu}(\lambda\cdot)=\tilde{\nu}$ for $\lambda>0$. It is enough to check that for any $\lambda,R>0$, $\tilde{\nu}(\lambda\cdot)\mathds{1}_{\cdot\subset R\mathbb{D}}=\tilde{\nu}(\cdot)\mathds{1}_{\cdot\subset R\mathbb{D}}$. This equality is a consequence of the fact that $\tilde{\nu}$ is a weak restriction measure and the scale invariance of $\Lambda^{*}$, proving the desired claim. Now, scale invariance of $\tilde{\nu}$ implies that (up to rescaling by a multiplicative constant) $\tilde{\nu}=\pi^{-1/2}\mathrm{d}c\otimes\tilde{\nu}^{\#}$ for some probability measure $\tilde{\nu}^{\#}$ on $\mathcal{E}$. Let $\tilde{\mathbb{E}}^{\#}$ be the expectation w.r.t. $\tilde{\nu}^{\#}$. At this stage, we do not know that $\nu$ is invariant under $z\mapsto 1/\bar{z}$. So we do not know yet that $\Theta$ is self-adjoint w.r.t. $L^{2}(\tilde{\nu})$ and that $\tilde{\mathcal{Q}}(F,G):=\langle F,\Theta G\rangle_{L^{2}(\tilde{\nu})}$ is a Hermitian form on $L^{2}(\tilde{\nu})$. This will be a consequence of the uniqueness we are currently proving.
By definition, the family of measures $\mathrm{d}\tilde{\nu}_{D}(\eta):=\mathds{1}_{\eta\subset D}e^{\frac{c_{\mathrm{m}}}{2}\Lambda^{*}(\eta,\partial D)}\mathrm{d}\tilde{\nu}(\eta)$ indexed by simply connected domains $D\subset\mathbb{C}$ containing 00 is conformally invariant. Now, the conclusion of Proposition 4.6 remains true if we replace $\nu_{\mathbb{D}}$ by $\tilde{\nu}_{\mathbb{D}}$ (the proof only uses this conformal restriction property), so $\tilde{\nu}_{\mathbb{D}}$ satisfies the integration by parts formula of Proposition 4.6. To get an integration by parts in the sphere, we proceed as before by noting that $|\log\epsilon|^{\frac{c_{\mathrm{m}}}{2}}\tilde{\nu}_{\epsilon\mathbb{D}^{*}}\to\tilde{\nu}$ weakly as $\epsilon\to 0$. Indeed, by the restriction property of $\tilde{\nu}$,
@@
and $\Lambda^{*}(\eta,\epsilon\mathbb{S}^{1})=-\log|\log\epsilon|+o(1)$ (this term does not depend on the specific choice of weak restriction measure on $\mathcal{J}_{0,\infty}$ and is thus the same as the one in [Zha21, page 385]). It follows that $\tilde{\nu}$ also satisfies the integration by parts formula of Theorem 4.1, i.e. we have the relations $\mathbf{L}_{n}^{*}=\Theta\circ\mathbf{L}_{-n}\circ\Theta$ and $\bar{\mathbf{L}}_{n}^{*}=\Theta\circ\bar{\mathbf{L}}_{-n}\circ\Theta$ for all $n\in\mathbb{Z}$, where the adjoint is w.r.t. to the $L^{2}(\tilde{\nu})$-inner product. In particular, $\tilde{\mathcal{Q}}$ satisfies the conclusions of Theorem 1.7 with $\tilde{\mathrm{R}}(\lambda):=\tilde{\mathbb{E}}^{\#}[e^{-2\lambda\tilde{c}_{0}}]$ in place of $\mathrm{R}(\lambda)$.
@@
7. Applications
We now discuss two consequences of the characterisation of the SLE loop measure in terms of restriction property.
7.1. Reversibility
The reversibility of the SLE loop measure is its invariance under the map $\iota:z\mapsto 1/\bar{z}$, which is present in Zhan’s axiomatisation. This property is not obvious a priori: in the chordal case for instance, Schramm’s definition of SLE only uses scale invariance [Sch00], but not the full Möbius invariance. In fact, the key property of invariance under inversion $z\mapsto 1/z$ was not part of the axioms and remained to be proved. This proof came only several years later, in [Zha08b] for the simple case $\kappa\in(0,4]$, and in [MS16c] for the non-simple case $\kappa\in(4,8)$.
Now, we explain how Theorem 1.9 provides an alternative and simpler approach to the question of reversibility. We remind the reader that our proof of uniqueness uses only the scale invariance of the SLE loop measure (which allows to write $\nu=\mathrm{d}c\otimes\nu^{\#}$), and not the full Möbius invariance (hence the argument is not recursive). By invariance of the Brownian loop measure under complex conjugation and Möbius inversion, it follows immediately from Definition 1.8 that $\iota^{*}(\mathrm{d}c\otimes\nu^{\#})$ is also a restriction measure, hence $\iota^{*}(\mathrm{d}c\otimes\nu^{\#})=K\mathrm{d}c\otimes\nu^{\#}$ for some $K>0$. To see that $K=1$, observe that for all $F\in\mathcal{S}_{\mathrm{exp}}(\mathbb{R})$ with $\int_{\mathbb{R}}F(c)\mathrm{d}c=1$, we have
$1=\int_{\mathbb{R}}F(c)\mathrm{d}c=K\mathbb{E}^{\#}[\int F(\tilde{c})\mathrm{d}c]=K\mathbb{E}^{\#}[\int_{\mathbb{R}}F(c-\tilde{c}_{0})\mathrm{d}c]=K\mathbb{E}^{\#}[\int_{\mathbb{R}}F(c)\mathrm{d}c]=K\int_{\mathbb{R}}F(c)\mathrm{d}c=K.$
As mentioned to us by W. Werner, and before the proof of Zhan [Zha08b] (see also [MS16b] and [LY21]), such an approach for proving reversibility of SLE had been considered, but a characterisation of SLE in terms of restriction property was missing; see for instance [Wer05, Section 8.1]. Theorem 1.9 is this missing piece, at least in the loop context.
7.2. Duality
Our second application concerns the duality $\kappa\leftrightarrow 16/\kappa$ for SLE. The duality of chordal SLE was first conjectured by Duplantier (see also Dubédat [Dub05]) and later proved by Zhan [Zha08a], Dubédat [Dub09a] and Miller–Sheffield [MS16a]. Duality for the SLE loop, as stated in Theorem 7.1 below, appears to be new. We mention that the paper in preparation [ACSW] gives an independent proof of this result, based on conformal welding of quantum discs and Lévy excursion theory (private communication with the authors).
[…]
Theorem 7.1 (Duality).
[…]
Theorem 7.1 does not address the self-dual case $\kappa^{\prime}=\kappa=4$. In this case, the curve is almost surely simple, and the duality for SLE4 can be understood as the reversibility proved in Section 7.1.
By conformal restriction, Theorem 7.1 implies a similar result in any simply connected domain $D$, where one looks at the boundary of the connected component which contains the boundary $\partial D$.
Proof.
[…]
$\mathds{1}_{\eta^{\prime}\subset D}=\mathds{1}_{\eta\subset D}\quad\text{and}\quad\Lambda^{*}(\eta^{\prime},D^{c})=\Lambda^{*}(\eta,D^{c}).$
[…]
We have just proved that the shape measure of the outer boundary of $\eta^{\prime}$ is the SLEκ shape measure $\nu_{\kappa}^{\#}$. This implies that $\mu_{\kappa}=K\nu_{\kappa}$ for some global multiplicative constant $K>0$. Our choice of normalisation ensures that the distribution of the $\log$-conformal radius of $\eta$ viewed from $\infty$ is $\frac{\mathrm{d}\tilde{c}}{\sqrt{\pi}}$. This coincides with our choice of normalisation for $\nu_{\kappa}$, so indeed $K=1$. ∎
@@ 8.7. Couplings with the Gaussian free field
Following Sheffield’s seminal article [She16], many couplings have been proved between SLE and the Gaussian free field (GFF). This includes a recent proof that the independent pair of an SLE loop with a quantum sphere (roughly speaking, the GFF on the sphere) can be obtained as the conformal welding of two independent quantum discs conditioned on their quantum length [AHS24].